OpenStudy (nicole143):

Can some one please explain how to write the equation of a parabola in vertex form? Axis of symmetry : x = -7 max height of 4 and also passes through the point (-6,0)

OpenStudy (nicole143):

Please don't just give the answer. I really need to know how.

OpenStudy (ddcamp):

Axis of symetry = x value of vertex (= h) max/min height = y value of vertex (= k) Now you have the equation: \[y = a(x-h)^2 + k\] We know h and k, to find a, plug in the given values of x and y (the point on the parabola)

OpenStudy (nicole143):

Okay, please don't go, I'm going to try it.

OpenStudy (nicole143):

A = 4?

OpenStudy (nicole143):

@DDCamp

OpenStudy (nicole143):

Would the equation be y = 4 (x + 6)^2 + 0 ??

OpenStudy (ddcamp):

We know h=-7, and k=4 (from the axis of symmetry and the max height) We know that the point (-6,0) is on the graph (x = -6 when y=0) \[y = a(x+7)^2+4 \\ 0 = a(-6+7)^2 + 4\]

OpenStudy (nicole143):

Oh, I did them the other way. so a = -4?

OpenStudy (nicole143):

And the equation would be y = -4(x + 7)^2 + 4 ??

OpenStudy (ddcamp):

Yup

OpenStudy (nicole143):

Is that the final equation/answer? there's no factoring or anything of the nature?

OpenStudy (ddcamp):

Nope. That's in vertex form.

OpenStudy (nicole143):

Thank you so much! @DDCamp